On a multiple choice quiz with 24 questions, each has 5 possible answers, find the variance of the number of correct answers.
Please help me set this up with the numbers provided. Not just the formula, it's the formula that is confusing me!
Thanks!

Answer by stanbon(75887)

On a multiple choice quiz with 24 questions, each has 5 possible answers, find the variance of the number of correct answers.
-----
Binomial Problem with n = 24 and p(correct) = 1/5
----
variance = sqrt(npq) = sqrt(24*1/5*4/5) = (4/5)sqrt(6)
====================
Cheers,
Stan H.
============